# Variations in Relevance Judgments and the Shelf Life of Test Collections

### v1

@@ 4.2. Pilot Study Findings
We conducted a pilot study over the top 10 pooled texts for 10 randomly sampled topics. Table 3 summarizes the results. We observe that the prevalence of grades 1 and 3 is inflated in these annotations compared to the original annotations. In discussions with annotators, there was ambiguity about what separates a relevant text from a highly relevant text. Additionally, cultural context leads to ambiguity in interpreting a query; for example, a query “dog day afternoon meaning” was difficult to interpret without broader context as a film exists with that name. The query could be interpreted as the meaning of the phrase or the film. From these ambiguities, we allowed annotators to use a controlled search engine to familiarise themselves with the topic. Annotators not from the USA generally noted the difficulty in USA-centric topics.
@@ 4.3. Core Assessment of Agreement
Quantification of relevance is confounded by many factors with several works reporting what is generally considered low agreement among annotators (Lesk and Salton, 1968; Burgin, 1992; Voorhees, 2000). Voorhees (2000) report overlap among annotators as a measure of agreement observing values around $0.4$. In Table 4, we present agreement measurements over both 4 grade and binarized relevance. We observe that under 4-grade relevance, all agreement values degrade relative to numbers stated in previous studies. Fleiss’ $\kappa$ measured over primary and secondary annotators indicates near-random agreement but generally improves by around 10 points in all cases over binary judgments. Overlap values over binary judgments generally correspond to those observed by Voorhees (2000), and over solely secondary annotators, we observe high overlap values, suggesting that while disagreement stems from the subjectivity of relevance and the annotation setting, the process of annotating over 4 grades is generally more variable regardless of topic granularity. In the third group, we find the lowest values in both 4 grade and binary relevance, which is again reflected in low overlap and Cohen’s $\kappa$ values, potentially suggesting diverging query intents or that this allocation of queries was particularly ambiguous. However, we see increases in overlap values when measured solely over secondary annotators, again suggesting that the annotation setting under which judgments are collected contributes to the improved agreement, as noted by Cuadra and Katter (1967) and Regazzi (1988) in previous annotation studies. Values of $\kappa$ are low in all cases, suggesting that generally quantifying relevance over 4 grades is more challenging for annotators than the re-annotation setting of Voorhees (2000). The lack of pre-defined narratives clarifying intent means annotators could interpret a query as they please, potentially resulting in three diverging topic interpretations.
@@
A figure showing the flow of grades from primary to secondary annotators, the key observations are that judgements generally reduce e.g. 1 to 0, 2 to 1..
In Figure 1, observe a Sankey flow diagram between the original relevance grade assigned by primary annotators and the new grades by secondary annotators. Like previous work, a secondary annotation setting often leads to the reduction of relevance grades, e.g., 1 to 0 or 2 to 1. Validating our decision to sample a subset of non-relevant judgements, 13% of 800 non-relevant texts change grade, with 77% being a change to a grade of 1; we analyse outlier cases (providing a grade of 2 or more) in Section 4.4. Measuring frequencies agnostic of previous judgments, grades 1 and 2 are annotated similarly across annotators, while grades 0 and 3 were annotated with greater variance. This variance in the case of grade 0 is generally consistent across annotators, and we provide a deeper analysis of grade 3 due to varying notions of relevance below.
@@ 4.4. Narratives, Intent, and Relevance
The observations of grade transitions in our work and generally low agreement noted by prior work, even in a binary setting, may be attributed to several factors noted in Section 2. Though we cannot fully isolate factors in the subjectivity of relevance, we choose to investigate the effect of narratives on inter-annotator agreement, given that recent corpora often do not have fixed narratives (Craswell et al., 2019, 2020), unlike those used in prior re-annotation studies (Harman, 1995). We select topics by median, minimum and max Fleiss’ $\kappa$ values from the annotation pool and consider where disagreement occurs and how narratives differ. Chosen examples illustrate different aspects of agreement, differing query intents, and a narrative’s specificity.
@@ 4.5. Narrative and Intent Ambiguity
Given the lack of public narratives provided in the release of TREC Deep Learning 2019, it is infeasible to isolate query intent as a factor in reliability without first collecting narratives. In fixing narratives, we produce an additional set of relevance judgements for queries of low, median, and high agreement by Fleiss $\kappa$, including the primary annotator. In Figure 2, We classify these queries as difficult, median and easy with respect to query intent determination and measure Cohen’s $\kappa$ between the judgements of both narrative groupings. It is clear that a fixed narrative reduces ambiguity with Cohen’s $\kappa$, improving by 24% and 46% for each annotation group between the difficult and median query; however, between the median and easy query, a decrease of 8% and an increase of 19% is observed.
@@ 4.6. Core Assessment of System Order
Figure 3. Comparing the stability of judgements through nDCG@10 and system swap probabilities over 10000 judgement permutations. We filter lexical comparisons to improve visibility in the central mass. ‘LLM-Neural’ denotes that a pair contains one LLM-based and one Neural-based ranker.
@@
Shows that LLM and Neural comparisons are frequently unstable under permutations.
@@
Shows that while some more traditional neural systems remain stable, depending on model capacity and distillation source overfitting can occur to the original DL’19 judgements.
@@
The core objective in re-annotation studies is to validate that downstream evaluation is not compromised by human bias in subjective relevance judgements. By measuring the stability of system order and the correlation of system order by target metrics, we can assess collection reliability. Following Voorhees (2000) and as described in Section 3.1, we sample permutations of relevance judgements and observe the stability of pair-wise system comparisons in Figure 3. Comparisons closest to the top right of the figure represent those where nDCG differences are large; however, they frequently swap in system order. A clear pareto-frontier is composed of comparisons between LLM and neural systems under judgement permutations. Lexical-neural comparisons generally lie in the main body of neural-neural and llm-llm comparisons, suggesting similar stability as was noted by Voorhees (2000) over TREC-6. However, more generally, we replicate the findings of previous works that despite low annotator agreement, under re-annotation, system ordering is highly correlated with the original DL’19 annotations even under 4-grade relevance. In Table 5, observe aggregate correlation values using both natural combinations of annotators and the in-sample approach described by Voorhees (2000).
@@
As the In-Sample approach includes primary judgements, a marginally higher correlation compared to combinations of solely secondary annotators is found. To investigate where system order diverges, we analyse ranking changes at a system level. In Figure 4 see that generally, new annotations led to a reduction in absolute nDCG@10 values relative to the original annotations, effectiveness measures on original annotations order systems, and that system order is not fully stable but LLM-based or distilled systems appear to be outliers, as such we explicitly measure the changes in system order concentrated around these systems. In Figure 5, we combine all possible permutations of annotators across annotation pairs. We measure changes in system order under these combinations and aggregate by system. Many lexical systems remain stable in ordering ($\Delta\approx 0$). Older neural approaches in the original system pool display higher variance but can improve or degrade in rank without significant difference. When considering only modern systems, we observe that smaller models trained with supervised and semi-supervised approaches consistently improve system rank. Within this group, however, larger semi-supervised models are more robust than their smaller counterparts under new judgments. List-wise encoders generally degrade in rank. These models use an interaction token to facilitate inter-document attention, assisting in out-of-domain effectiveness (Schlatt et al., 2024a). Changes in effectiveness are more pronounced and are significant in smaller distilled language models. This is notable as though we see continuing improvement in smaller models out-of-domain through distillation. One should be cautious that in-domain effectiveness may be less robust than benchmarks indicate, which for many applications may be more important than robust effectiveness across heterogeneous corpora.
@@ 4.7. Systems versus Humans
Recall our argument that a “perfect” score on a normalized metric may not be a feasible or desirable goal in ranking evaluation; as such, we measure the effectiveness of human annotators aggregated in multiple forms outlined in Section 3.3. In Table 6, see that systems can exceed or reach parity with secondary annotators, particularly with respect to nDCG@10 and MRR@10, with multiple systems, both LLM and encoder-based exceeding annotators. When excluding unjudged documents (which is known to overestimate the retrieval effectiveness (Fröbe et al., 2023; Sakai, 2008)), annotator performance measured by precision can be exceeded by neural systems across nDCG, P, and MRR; however, recall remains lower in all settings. Given our findings in Figure 4, an observation that systems are better than humans should be considered cautiously. However, the robustness of the LLM-based RankZephyr observed both by system order robustness over new judgments and its parity with human annotators suggests that with sufficient capacity, a smaller system can be distilled from larger systems while remaining robust. Observing similar effectiveness to humans and confidence intervals providing no way to discriminate systems, we consider that achieving significant improvements would suggest far exceeding humans, which may not be possible in a way conducive to improving information access and may solely serve to improve precision on specific topics and intents. The ability of distilled systems to reach parity with human effectiveness across multiple annotator combinations while obtaining largely reduced recall in a judged setting, even when re-ranking a neural first-stage retriever, is undesirable for several downstream tasks. Although precision is approaching or overtaking the effectiveness of an oracle, recall remains a challenge in densely annotated collections.
@@ 5. Conclusion
We considered important studies on the impact of the annotator agreement on retrieval evaluations in light of modern neural evaluation scenarios. In reproducing re-annotation processes and evaluations by Voorhees (2000) on a modern test collection, we validate the stability system ordering under re-annotation under 4-grade relevance and ambiguous query intent. We raised concerns that the broad and frequent use of test collections to “tune” learned systems implicitly through influences from published works may reduce the reliability of a collection over time. We identified that the subjectivity of relevance through query intent may be a factor in determining “overfitting” in ranking tasks. As a query may have several interpretations, we re-annotate the popular TREC DL’19 collection. We found that system order varies when evaluating large language models and systems distilled from them. Furthermore, we observed that the current state-of-the-art can outperform combinations of human annotators on original relevance judgments, suggesting we may have reached a realistic bound on precision for this collection. We posit that further improvement on this collection may not indicate that a system is better than another in a meaningful way. Our process has limitations—most notably, re-annotation is incredibly costly. Future work will investigate how to draw similar conclusions about a collection’s expiration without the need for extensive human effort.
@@ Acknowledgements.
We thank Sean MacAvaney and others involved in making the sessions of the Collab-a-thon at ECIR‘24 (MacAvaney et al., 2024) an engaging space for collaboration.

### v2

@@ 4.2. Pilot Study Findings
We conducted a pilot study over the top 10 pooled texts for 10 randomly sampled topics. Table 3 summarizes the results. We observe that the prevalence of grades 1 and 3 is inflated in these annotations compared to the original annotations. In discussions with annotators, there was ambiguity about what separates a relevant text from a highly relevant text. Additionally, cultural context leads to ambiguity in interpreting a query; for example, a query “dog day afternoon meaning” was difficult to interpret without broader context, as a film exists with that name. The query could be interpreted as the meaning of the phrase or the film. From these ambiguities, we allowed annotators to use a controlled search engine to familiarise themselves with the topic. Annotators not from the USA generally noted the difficulty in USA-centric topics; however, we leave bias and query difficulty by nationality to future work, as this line of study is not central to our hypotheses.
@@ 4.3. Core Assessment of Agreement
Quantification of relevance is confounded by many factors with several works reporting what is generally considered low agreement among annotators (Lesk and Salton, 1968; Burgin, 1992; Voorhees, 2000). Voorhees (2000) report overlap among annotators as a measure of agreement, observing values around $0.4$. In Table 4, we present agreement measurements over both 4-grade and binarized relevance. We observe that under 4-grade relevance, all agreement values degrade relative to those stated in previous studies. Fleiss’ $\kappa$ measured over primary and secondary annotators indicates near-random agreement but generally improves by around 10 points in all cases over binary judgments. Overlap values over binary judgments generally correspond to those observed by Voorhees (2000), and over solely secondary annotators, we observe high overlap values, suggesting that while disagreement stems from the subjectivity of relevance and the annotation setting, the process of annotating over 4 grades is generally more variable regardless of topic granularity. In the third group, we find the lowest values in both 4 grade and binary relevance, which is again reflected in low overlap and Cohen’s $\kappa$ values, potentially suggesting diverging query intents or that this allocation of queries was particularly ambiguous. However, we see increases in overlap values when measured solely over secondary annotators, again suggesting that the annotation setting under which judgments are collected contributes to the improved agreement, as noted by Cuadra and Katter (1967) and Regazzi (1988) in previous annotation studies. Values of $\kappa$ are low in all cases, suggesting that generally quantifying relevance over 4 grades is more challenging for annotators than the re-annotation setting of Voorhees (2000). The lack of pre-defined narratives clarifying intent means annotators could interpret a query as they please, potentially resulting in three diverging topic interpretations.
@@
A figure showing the flow of grades from primary to secondary annotators, the key observations are that judgments generally reduce e.g. 1 to 0, 2 to 1..
In Figure 1, observe a Sankey flow diagram between the original relevance grade assigned by primary annotators and the new grades by secondary annotators. Like previous work, a secondary annotation setting often leads to the reduction of relevance grades, e.g., 1 to 0 or 2 to 1. Validating our decision to sample a subset of non-relevant judgments, 13% of 800 non-relevant texts change grade, with 77% being a change to a grade of 1; we analyse outlier cases (providing a grade of 2 or more) in Section 4.4. Measuring frequencies agnostic of previous judgments, grades 1 and 2 are annotated similarly across annotators, while grades 0 and 3 were annotated with greater variance. This variance in the case of grade 0 is generally consistent across annotators, and we provide a deeper analysis of grade 3 due to varying notions of relevance below.
@@ 4.4. Narratives, Intent, and Relevance
The observations of grade transitions in our work and generally low agreement noted by prior work, even in a binary setting, may be attributed to several factors noted in Section 2. Though we cannot fully isolate factors in the subjectivity of relevance, we choose to investigate the effect of narratives on inter-annotator agreement, given that recent corpora often do not have fixed narratives (Craswell et al., 2019, 2020), unlike those used in prior re-annotation studies (Harman, 1995). We select topics by median, minimum, and max Fleiss’ $\kappa$ values from the annotation pool and consider where disagreement occurs and how narratives differ. Chosen examples illustrate different aspects of agreement, differing query intents, and a narrative’s specificity.
@@ 4.5. Narrative and Intent Ambiguity
Given the lack of public narratives provided in the release of TREC Deep Learning 2019, it is infeasible to isolate query intent as a factor in reliability without first collecting narratives. In fixing narratives, we produce an additional set of relevance judgments for queries of low, median, and high agreement by Fleiss $\kappa$, including the primary annotator. In Figure 2, We classify these queries as difficult, median and easy with respect to query intent determination and measure Cohen’s $\kappa$ between the judgments of both narrative groupings. It is clear that a fixed narrative reduces ambiguity with Cohen’s $\kappa$, improving by 24% and 46% for each annotation group between the difficult and median query; however, between the median and easy query, a decrease of 8% and an increase of 19% is observed.
@@ 4.6. Core Assessment of System Order
Figure 3. Comparing the stability of judgments through nDCG@10 and system swap probabilities over 10000 judgment combinations. We filter lexical comparisons to improve visibility in the central mass. ‘LLM-Neural’ denotes that a pair contains one LLM-based and one Neural-based ranker.
@@
Shows that LLM and Neural comparisons are frequently unstable under combinations.
@@
Shows that while some more traditional neural systems remain stable, depending on model capacity and distillation source overfitting can occur to the original DL’19 judgments.
@@
The core objective in re-annotation studies is to validate that downstream evaluation is not compromised by human bias in subjective relevance judgments. By measuring the stability of system order and the correlation of system order by target metrics, we can assess collection reliability. Following Voorhees (2000) and as described in Section 3.1, we sample combinations of relevance judgments and observe the stability of pair-wise system comparisons in Figure 3. Comparisons closest to the top right of the figure represent those where nDCG differences are large; however, they frequently swap in system order. A clear pareto-frontier is composed of comparisons between LLM and neural systems under judgment combinations. Lexical-neural comparisons generally lie in the main body of neural-neural and llm-llm comparisons, suggesting similar stability as was noted by Voorhees (2000) over TREC-6. However, more generally, we replicate the findings of previous works that despite low annotator agreement, under re-annotation, system ordering is highly correlated with the original DL’19 annotations even under 4-grade relevance. In Table 5, observe aggregate correlation values using both natural combinations of annotators and the in-sample approach described by Voorhees (2000).
@@
As the In-Sample approach includes primary judgments, a marginally higher correlation compared to combinations of solely secondary annotators is found. To investigate where system order diverges, we analyse ranking changes at a system level. In Figure 4 see that generally, new annotations led to a reduction in absolute nDCG@10 values relative to the original annotations, effectiveness measures on original annotations order systems, and that system order is not fully stable but LLM-based or distilled systems appear to be outliers, as such we explicitly measure the changes in system order concentrated around these systems. In Figure 5, we combine all possible combinations of annotators across annotation pairs. We measure changes in system order under these combinations and aggregate by system. Many lexical systems remain stable in ordering ($\Delta\approx 0$). Older neural approaches in the original system pool display higher variance but can improve or degrade in rank without a significant difference. When considering only modern systems, we observe that smaller models trained with supervised and semi-supervised approaches consistently improve system rank. Within this group, however, larger semi-supervised models are more robust than their smaller counterparts under new judgments. List-wise encoders generally degrade in rank. These models use an interaction token to facilitate inter-document attention, assisting in out-of-domain effectiveness (Schlatt et al., 2024a). Changes in effectiveness are more pronounced and are significant in smaller distilled language models. This is notable as though we see continuing improvement in smaller models out-of-domain through distillation. One should be cautious that in-domain effectiveness may be less robust than benchmarks indicate, which for many applications may be more important than robust effectiveness across heterogeneous corpora.
@@ 4.7. Systems versus Humans
Recall our argument that a “perfect” score on a normalized metric may not be a feasible or desirable goal in ranking evaluation; as such, we measure the effectiveness of human annotators aggregated in multiple forms outlined in Section 3.3. In Table 6, see that systems can exceed or reach parity with secondary annotators, particularly with respect to nDCG@10 and MRR@10, with multiple systems, both LLM and encoder-based, exceeding annotators. When excluding unjudged documents (which is known to overestimate the retrieval effectiveness (Fröbe et al., 2023; Sakai, 2008)), annotator performance measured by precision can be exceeded by neural systems across nDCG, P, and MRR; however, recall remains lower in all settings. Given our findings in Figure 4, an observation that systems are better than humans should be considered cautiously. However, the robustness of the LLM-based RankZephyr observed both by system order robustness over new judgments and its parity with human annotators suggests that with sufficient capacity, a smaller system can be distilled from larger systems while remaining robust. Observing similar effectiveness to humans and confidence intervals providing no way to discriminate systems, we consider that achieving significant improvements would suggest far exceeding humans, which may not be possible in a way conducive to improving information access and may solely serve to improve precision on specific topics and intents. The ability of distilled systems to reach parity with human effectiveness across multiple annotator combinations while obtaining largely reduced recall in a judged setting, even when re-ranking a neural first-stage retriever, is undesirable for several downstream tasks. Although precision is approaching or overtaking the effectiveness of an oracle, recall remains a challenge in densely annotated collections.
@@ 5. Conclusion
We considered important studies on the impact of the annotator agreement on retrieval evaluations in light of modern neural evaluation scenarios. In reproducing re-annotation processes and evaluations by Voorhees (2000) on a modern test collection, we validate the stability of system ordering under re-annotation under 4-grade relevance and ambiguous query intent. We raised concerns that the broad and frequent use of test collections to “tune” learned systems implicitly through influences from published works may reduce the reliability of a collection over time. We identified that the subjectivity of relevance through query intent may be a factor in determining “overfitting” in ranking tasks. As a query may have several interpretations, we re-annotate the popular TREC DL’19 collection. We found that system order varies when evaluating large language models and systems distilled from them. Furthermore, we observed that the current state-of-the-art can outperform combinations of human annotators on original relevance judgments, suggesting we may have reached a realistic bound on precision for this collection. We posit that further improvement on this collection may not indicate that a system is better than another in a meaningful way. Our process has limitations—most notably, re-annotation is incredibly costly. Future work will investigate how to draw similar conclusions about a collection’s expiration without the need for extensive human effort.
@@ Acknowledgements.
We thank all involved in making the sessions of the Collab-a-thon at ECIR‘24 (MacAvaney et al., 2024) an engaging space for collaboration.